Please help solve!
Find the equation in standard form for each line that satifies the given conditions:
a) parallel to 5x + y = 3 and passes through (0,1)
b) parallel to the y axis and passes through (1, -2)
Thanks so much!

Answer by mathispowerful(115)

a) parallel to 5x + y = 3 and passes through (0,1)
b) parallel to the y axis and passes through (1, -2)

answer:
a) "parallel to 5x + y = 3" indicates they have same slope.
the slope in the given equation is -5.
so let y = -5x + b represent the equation in the question.
since it passes through (0,1), replace x by 0, y by 1, we get
1 = b, so the answer is y = -5x + 1
standard format: 5x + y = 1
b) since the line is parallel to the y axis, its x value won't change,
so the equation is
x = 1